OpenStudy (anonymous):

What is 30% in its simplest fraction

Miracrown (miracrown):

Answer: 3/10\[\frac{ 30% }{ 100% } = \frac{ 30 }{ 100 } = \frac{ 30 }{ 10 } / \frac{ 100 }{ 10} = \frac{ 3 }{ 10 }\]

mathslover (mathslover):

O.O She used LaTeX? @Miracrown ?